# Speed, Accuracy, and Complexity

### v1

@@ 6.  Identifying Complexity
However, in many situations, the analyst does not know which choices are optimal. For instance, suppose that one is interested in assessing which features of a particular mechanism that elicits private information make it more complex. In this case, which choices are optimal will typically depend exactly on the private information the designer is trying to elicit. Furthermore, our earlier results highlight the limitations of using expected response time as a measure of problem complexity.131313 We found that this non-monotonicity is present even if we consider other moments of the distribution, such as variance, even if the population is homogeneous in preferences and ability. Hence, we would like some behavioural measure that is monotone in complexity that can be easily observed, aggregated across heterogeneous populations, and used to infer the complexity of a problem.
@@ 7.  Final Remarks
We believe that our results have important implications for the use of response times in the social sciences. While response times provide an important source of information that is able to shed light on the cognitive processes underlying decision-making, our results support existing evidence that, on their own, they are not a reliable indicator of problem complexity or ability. Instead, we propose to use small incentive distortions to infer problem complexity. These are easy to implement in real-life settings and could be used to study which features make specific problems more or less complex, and evaluate if particular policies improve them. We find these to be promising avenues for future research.
@@ A.1.  Proof of Lemma 1
We focus on the $V(p)=\delta V_{a}(p)+u(b,p)$ and $\operatorname*{arg\,max}_{\tau}\mathbb{E}_{p_{0}}[v(p_{\tau})-c\tau]=\operatorname*{arg\,max}_{\tau}\mathbb{E}_{p_{0}}[v_{a}(p_{\tau})-(c/\delta)\tau]$, since the other case is analogous. Note that $u(\alpha,p)=\delta u_{a}(\alpha,p)+u(b,p)$, for $\alpha=a,b$. For any stopping time $\tau$ which is finite almost surely, $\mathbb{E}_{p_{0}}[v(p_{\tau})-c\tau]=\delta\mathbb{E}_{p_{0}}[(p_{\tau}-\tilde{p})^{+}-(c/\delta)\tau]+\mathbb{E}_{p_{0}}[u(b,p_{\tau})]=\delta\mathbb{E}_{p_{0}}[v_{a}(p_{\tau})-(c/\delta)\tau]+\mathbb{E}_{p_{0}}[u(b,p_{\tau})]=\delta\mathbb{E}_{p_{0}}[v_{a}(p_{\tau})-(c/\delta)\tau]+u(b,p_{0})$, where we used optional stopping theorem with the fact that beliefs are a martingale and lie in a compact set. Any optimal stopping time must be finite almost surely, as otherwise the expected value is negative infinity. Hence, $\operatorname*{arg\,max}_{\tau}\mathbb{E}_{p_{0}}[v(p_{\tau})-c\tau]=\operatorname*{arg\,max}_{\tau}\delta\mathbb{E}_{p_{0}}[v_{a}(p_{\tau})-(c/\delta)\tau]+u(b,p_{0})=\operatorname*{arg\,max}_{\tau}\mathbb{E}_{p_{0}}[v_{a}(p_{\tau})-(c/\delta)\tau]$ and $\max_{\tau}\mathbb{E}_{p_{0}}[v(p_{\tau})-c\tau]=\delta\max_{\tau}\mathbb{E}_{p_{0}}[v_{a}(p_{\tau})-(c/\delta)\tau]+u(b,p_{0})=\delta V_{a}(p_{0})+u(b,p_{0})$. ∎
@@ A.2.  Proof of Lemma 2
A.3.  Proof of Lemma 3
Proof.
We focus on $V_{a}$; the proof for $V_{b}$ is analogous. Let $g=(g_{t})$ denote a stochastic process adapted to the natural filtration induced by $(X_{t})$. Then, note that $V_{a}(p_{0})=\sup_{g,\tau}\mathbb{E}_{p_{0}}[u_{a}(g_{\tau},p_{\tau})-c\tau]$. Letting $W_{a}(p_{0},\tau,g):=\mathbb{E}_{p_{0}}[u_{a}(g_{\tau},p_{\tau})-c\tau]=\mathbb{E}_{p_{0}}[u_{a}(g_{\tau},\theta)-c\tau]$, we have that $V_{a}(p_{0})=\sup_{g,\tau}W_{a}(p_{0},\tau,g)$. Since $W_{a}$ is linear in $p_{0}$, $V_{a}$ is convex in $p_{0}$.
@@ A.4.  Proof of Lemma 4
From Lemma 2, there are $10<\underline{p}\leq\overline{p}<$ 1 such that $\tau^{*}:=\inf\{t\geq 0:p_{t}\notin(\underline{p},\overline{p})\}$. Given this, the fact that $V$ is the unique viscosity solution of the free-boundary problem follows from OksendalReikvam1998SSR. ∎
@@ A.5.  Proof of Proposition 1
Since $V_{a}^{\prime\prime}$ exists a.e., from CrandallIshiiLions1992BullAMS, the HJB will hold with equality whenever $V_{a}$ is twice differentiable, i.e. almost everywhere on $(\underline{p},\overline{p})$. Furthermore, $0\leq V_{a}^{\prime\prime}\leq(\mu/\sigma)^{-2}\tilde{c}\max\{(\underline{p}(1-\underline{p}))^{-2},(\overline{p}(1-\overline{p}))^{-2}\}$, and so $V_{a}^{\prime\prime}$ is bounded whenever it exists. Take any sequence $(p_{n})\subset(\underline{p},\overline{p})$ such that $p_{n}\to p^{*}\in(\underline{p},\overline{p})$. Then, there is a subsequence $(p_{m})$ such that $V_{a}^{\prime}(p_{m})\to V_{a,\infty}^{\prime}$, $V_{a}^{\prime\prime}(p_{m})\to V_{a,\infty}^{\prime\prime}$, and, by continuity, $0=-\tilde{c}+(\mu/\sigma)^{2}\left(p^{*}(1-p^{*})\right)^{2}V_{a,\infty}^{\prime\prime}$. Hence $(V_{a,\infty}^{\prime},V_{a,\infty}^{\prime\prime})$ belongs to both the super- and subjet of $V_{a}$ at $p^{*}$, implying that $V_{a}$ is twice continuously differentiable everywhere except possibly at $\{\underline{p},\overline{p}\}$.
@@ A.6.  Proof of Theorem 1
Letting $\ell(p):=\ln(p/(1-p))$, we first note that the information process $X_{t}$ induces a process for log-odds beliefs given by $\ell_{t}:=2\frac{\mu}{\sigma^{2}}X_{t}$. Second, we observe that, given $\theta$, $\exp(-(2s(\theta)\mu/\sigma^{2})X_{t})$ is a martingale and as $X_{\tau}$ is therefore bounded, we can apply the optional stopping theorem (EthierKurtz2005Book, Ch. 2, Theorem 2.3) to obtain $1=\exp(-(2s(\theta)\mu/\sigma^{2})X_{0})=\mathbb{E}[\exp(-(2s(\theta)\mu/\sigma^{2})X_{\tau})\mid\theta]=\mathbb{E}[\exp(-s(\theta)\ell_{\tau})\mid\theta]$.
@@ A.7.  Proof of Proposition 2
We first note that the optimal effort for agent $i$, $e_{i}^{*}$, is constant over time and equal to $e^{*}_{i}=\lambda_{i}e^{*}$, where $e^{*}$ is such that $e^{*}c^{\prime}(e^{*})=c(e^{*})$. Write $c_{i}(e_{i}):=c(e_{i}/\lambda_{i})$. The argument (already in MoscariniSmith2001Ecta) is that, on the continuation region, the HJB equation is given by
@@ Appendix B.  Discounting
Formally, the decision-maker chooses a stopping time $\tau$ and a stochastic process $(\alpha_{t})$ taking values in $A$, both adapted to the natural filtration induced by the stochastic process$(X_{t})$ in order to maximise their expected payoff, $\mathbb{E}_{p_{0}}\left[\exp(-r\tau)(1_{\alpha_{\tau}=\theta})\right]$, where $p_{t}$ denotes the posterior belief about $\theta$ given $X_{t}$, and $v(p):=\max\{p,1-p\}$. Denote their value function by
$V(p_{0})=\sup_{\tau}\mathbb{E}_{p_{0}}\left[\exp(-r\tau)(1_{\alpha_{\tau}=\theta})\right].$
@@
$\mathbb{E}_{p_{0}}[\exp(-r\tau^{*})(1_{\alpha_{\tau^{*}}=\theta})]=\mathbb{E}_{p_{0}}[\exp(-r\tau^{*})(1_{\alpha_{\tau^{*}}=\theta})|\theta=a]=\frac{1}{2}\exp\left(\frac{\overline{\ell}}{2}\right)\text{sech}\left(\frac{\overline{\ell}}{2}\;\kappa\right)$
where the last equality follows from BorodinSalminen2002Book and $\kappa:={(2r(\sigma/\mu)^{2}+1)}^{1/2}$.
Taking first-order conditions, we get that the optimal stopping boundary is given by ¯*$\overline{\ell}^{*}=2\,\text{tanh}^{-1}(1/\kappa)/\kappa.$ It is then immediate that accuracy, $\mathbb{P}(\alpha_{\tau^{*}}=\theta)=(1+\exp(-\overline{\ell}^{*}))^{-1}$, is increasing in $\mu/\sigma$. The expected stopping time fixing $\overline{\ell}=-\underline{\ell}$ is known and given by (3), that is, $\mathbb{E}[\tau]=\frac{\overline{\ell}\,\tanh\left(\overline{\ell}/2\right)}{2(\mu/\sigma)^{2}}.$ Substituting for the optimal stopping threshold $\overline{\ell}^{*}$, we have $\mathbb{E}[\tau^{*}]=\frac{\kappa^{2}-1}{\kappa r}\,\text{coth}^{-1}(\kappa)\,\text{tanh}\left(\text{coth}^{-1}(\kappa)/\kappa\right),$ which is inverse $U$-shaped in $\mu/\sigma$, with $\lim_{\mu/\sigma\downarrow 0}\mathbb{E}[\tau^{*}]=\lim_{\mu/\sigma\uparrow\infty}\mathbb{E}[\tau^{*}]=0$ — see Figure 5.
Refer to caption (a) Refer to caption (b) Refer to caption (c)

### v2

@@ 6.  Identifying Complexity
However, in many situations, the analyst does not know which choices are optimal. For instance, suppose that one is interested in assessing which features of a particular mechanism that elicits private information make it more complex. In this case, which choices are optimal will typically depend exactly on the private information the designer is trying to elicit. Furthermore, our earlier results highlight the limitations of using expected response time as a measure of problem complexity.191919 We found that this non-monotonicity is present even if we consider other moments of the distribution, such as variance, even if the population is homogeneous in preferences and ability. Hence, we would like some behavioural measure that is monotone in complexity that can be easily observed, aggregated across heterogeneous populations, and used to infer the complexity of a problem.
@@ 7.  Final Remarks
We believe that our results have important implications for the use of response times in the social sciences. While response times provide an important source of information that is able to shed light on the cognitive processes underlying decision-making, our results support existing evidence that, on their own, they are not a reliable indicator of problem complexity or ability. Instead, we propose to use small incentive distortions to infer problem complexity. These are easy to implement in real-life settings and could be used to study which features make specific problems more or less complex, and evaluate if particular policies improve them. In current ongoing, Gonçalves et al. (2024) not only corroborate the results in Theorem 1 — already evidenced in the above-referenced literature — but also demonstrate that incentive manipulation can indeed be used to reveal problem complexity.
Our results also highlight the importance of recognising how complex a problem is, as well as one’s own ability in tackling problems of a particular nature. Several questions emerge naturally. Can we infer problem complexity via incentive manipulation also when the decision-maker is not immediately aware of the problem’s complexity? Would this kind of uncertainty, leading to individuals exerting too much or too little effort, be able to rationalise phenomena akin to over- and underconfidence? Are agents with different abilities able to learn to self-sort into tackling problems (activities or jobs) regarding which they hold a comparative advantage? We find these to be promising avenues for future research.
@@ A.1.  Proof of Lemma 1
We focus on the $V(p)=\delta V_{a}(p)+u(b,p)$ and $\operatorname*{arg\,max}_{\tau}\mathbb{E}[v(p_{\tau})-c\tau]=\operatorname*{arg\,max}_{\tau}\mathbb{E}[v_{a}(p_{\tau})-(c/\delta)\tau]$, since the other case is analogous. Note that $u(\alpha,p)=\delta u_{a}(\alpha,p)+u(b,p)$, for $\alpha=a,b$. For any stopping time $\tau$ which is finite almost surely, $\mathbb{E}[v(p_{\tau})-c\tau]=\delta\mathbb{E}[(p_{\tau}-\tilde{p})^{+}-(c/\delta)\tau]+\mathbb{E}[u(b,p_{\tau})]=\delta\mathbb{E}[v_{a}(p_{\tau})-(c/\delta)\tau]+\mathbb{E}[u(b,p_{\tau})]=\delta\mathbb{E}[v_{a}(p_{\tau})-(c/\delta)\tau]+u(b,p_{0})$, where we used optional stopping theorem with the fact that beliefs are a martingale and lie in a compact set. Any optimal stopping time must be finite almost surely, as otherwise the expected value is negative infinity. Hence, $\operatorname*{arg\,max}_{\tau}\mathbb{E}[v(p_{\tau})-c\tau]=\operatorname*{arg\,max}_{\tau}\delta\mathbb{E}[v_{a}(p_{\tau})-(c/\delta)\tau]+u(b,p_{0})=\operatorname*{arg\,max}_{\tau}\mathbb{E}[v_{a}(p_{\tau})-(c/\delta)\tau]$ and $\max_{\tau}\mathbb{E}[v(p_{\tau})-c\tau]=\delta\max_{\tau}\mathbb{E}[v_{a}(p_{\tau})-(c/\delta)\tau]+u(b,p_{0})=\delta V_{a}(p_{0})+u(b,p_{0})$. ∎
@@ A.2.  Proof of Lemma 2
Proof.
We focus on $V_{a}$; the proof for $V_{b}$ is analogous. Let $g=(g_{t})$ denote a stochastic process adapted to the natural filtration induced by ${(X_{t})}_{t}$. Note that $V_{a}(p_{0})=\sup_{g,\tau}\mathbb{E}[u_{a}(g_{\tau},p_{\tau})-c\tau]$. Letting $W_{a}(p_{0},\tau,g):=\mathbb{E}[u_{a}(g_{\tau},p_{\tau})-c\tau]=\mathbb{E}[u_{a}(g_{\tau},\theta)-c\tau]$, we have $V_{a}(p_{0})=\sup_{g,\tau}W_{a}(p_{0},\tau,g)$. Since $W_{a}$ is linear in $p_{0}$, $V_{a}$ is convex in $p_{0}$.
A.3.  Proof of Lemma 3
@@ A.4.  Proof of Lemma 4
From Lemma 3, there are $10<\underline{p}\leq\overline{p}<$ 1 such that $\tau^{*}:=\inf\{t\geq 0:p_{t}\notin(\underline{p},\overline{p})\}$. Given this, the fact that $V$ is the unique viscosity solution of the free-boundary problem follows from Øksendal and Reikvam (1998, Proposition 3.1.). ∎
@@ A.5.  Proof of Proposition 1
Since $V_{a}^{\prime\prime}$ exists a.e., from Crandall et al. (1992, p. 15), the HJB will hold with equality whenever $V_{a}$ is twice differentiable, i.e. almost everywhere on $(\underline{p},\overline{p})$. Furthermore, $0\leq V_{a}^{\prime\prime}\leq(\mu/\sigma)^{-2}\tilde{c}\max\{(\underline{p}(1-\underline{p}))^{-2},(\overline{p}(1-\overline{p}))^{-2}\}$, and so $V_{a}^{\prime\prime}$ is bounded whenever it exists. Take any sequence $(p_{n})\subset(\underline{p},\overline{p})$ such that $p_{n}\to p^{*}\in(\underline{p},\overline{p})$. Then, there is a subsequence $(p_{m})$ such that $V_{a}^{\prime}(p_{m})\to V_{a,\infty}^{\prime}$, $V_{a}^{\prime\prime}(p_{m})\to V_{a,\infty}^{\prime\prime}$, and, by continuity, $0=-\tilde{c}+(\mu/\sigma)^{2}\left(p^{*}(1-p^{*})\right)^{2}V_{a,\infty}^{\prime\prime}$. Hence $(V_{a,\infty}^{\prime},V_{a,\infty}^{\prime\prime})$ belongs to both the super- and subjet of $V_{a}$ at $p^{*}$, implying that $V_{a}$ is twice continuously differentiable everywhere except possibly at $\{\underline{p},\overline{p}\}$.
@@ A.6.  Proof of Theorem 1
Letting $\ell(p):=\ln(p/(1-p))$, we first note that the information process $X_{t}$ induces a process for log-odds beliefs given by $\ell_{t}:=2\frac{\mu}{\sigma^{2}}X_{t}$. Second, we observe that, given $\theta$, $\exp(-(2s(\theta)\mu/\sigma^{2})X_{t})$ is a martingale and as $X_{\tau}$ is therefore bounded, we can apply the optional stopping theorem (Ethier and Kurtz, 2005, Ch. 2, Theorem 2.3) to obtain $1=\exp(-(2s(\theta)\mu/\sigma^{2})X_{0})=\mathbb{E}[\exp(-(2s(\theta)\mu/\sigma^{2})X_{\tau})\mid\theta]=\mathbb{E}[\exp(-s(\theta)\ell_{\tau})\mid\theta]$.
@@ A.7.  Proof of Proposition 2
We first note that the optimal effort for agent $i$, $e_{i}^{*}$, is constant over time and equal to $e^{*}_{i}=\lambda_{i}e^{*}$, where $e^{*}$ is such that $e^{*}c^{\prime}(e^{*})=c(e^{*})$. Write $c_{i}(e_{i}):=c(e_{i}/\lambda_{i})$. The argument (already in Moscarini and Smith (2001)) is that, on the continuation region, the HJB equation is given by
@@ Appendix B.  Discounting
Formally, the decision-maker chooses a stopping time $\tau$ and a stochastic process $(\alpha_{t})$ taking values in $A$, both adapted to the natural filtration induced by the stochastic process ${(X_{t})}_{t}$ in order to maximise their expected payoff, $\mathbb{E}\left[\exp(-r\tau)(1_{\alpha_{\tau}=\theta})\right]$, where $p_{t}$ denotes the posterior belief about $\theta$ given $X_{t}$, and $v(p):=\max\{p,1-p\}$. Denote their value function by
$V(p_{0})=\sup_{\tau}\mathbb{E}\left[\exp(-r\tau)(1_{\alpha_{\tau}=\theta})\right].$
@@
$\mathbb{E}[\exp(-r\tau^{*})(1_{\alpha_{\tau^{*}}=\theta})]=\mathbb{E}[\exp(-r\tau^{*})(1_{\alpha_{\tau^{*}}=\theta})|\theta=a]=\frac{1}{2}\exp\left(\frac{\overline{\ell}}{2}\right)\text{sech}\left(\frac{\overline{\ell}}{2}\;\kappa\right)$
where the last equality follows from Borodin and Salminen (2002, p. 309, 3.0.5(b)) and $\kappa:={(2r(\sigma/\mu)^{2}+1)}^{1/2}$.
Taking first-order conditions, we get that the optimal stopping boundary is given by ¯$\overline{\ell}^{*}=2\,\text{tanh}^{-1}(1/\kappa)/\kappa.$ It is then immediate that accuracy, $\mathbb{P}(\alpha_{\tau^{*}}=\theta)=(1+\exp(-\overline{\ell}^{*}))^{-1}$, is increasing in $\mu/\sigma$. The expected stopping time fixing $\overline{\ell}=-\underline{\ell}$ is known and given by (3), that is, $\mathbb{E}[\tau]=\frac{\overline{\ell}\,\tanh\left(\overline{\ell}/2\right)}{2(\mu/\sigma)^{2}}.$ Substituting for the optimal stopping threshold $\overline{\ell}^{*}$, we have $\mathbb{E}[\tau^{*}]=\frac{\kappa^{2}-1}{\kappa r}\,\text{coth}^{-1}(\kappa)\,\text{tanh}\left(\text{coth}^{-1}(\kappa)/\kappa\right),$ which is inverse $U$-shaped in $\mu/\sigma$, with $\lim_{\mu/\sigma\downarrow 0}\mathbb{E}[\tau^{*}]=\lim_{\mu/\sigma\uparrow\infty}\mathbb{E}[\tau^{*}]=0$ — see Figure 5.